john needs to rent a theatre for a concert. The theatre charge $700 plus $3per person. john plans to selltickets for $10 per person. let C represent the cost of the eventand N the number of people attending the concert.
Write a linear system to represent the situation.
How many tickets must be sold to break evan?
what will be the total cost be?

Answer by Theo(13342)

c = 700 + 3 * n

c = cost
n = number of persons attending.

r = 10 * n

r = revenue
n = number of persons attending.

he will break even when r = c

this makes:

10 * n = 700 + 3 * n

subtract 3 * n from both sides of this equation to get:

7 * n = 700

divide both sides of this equation by 7 to get:

n = 100

he should break even when 100 peopled attend.

cost for 100 people = 700 + 300 = 1000

revenue for 100 people = 10 * 100 = 1000

that's the break even point.